Given the balanced equation, calculate the mass of product that can be prepared from 3.00 g of zinc metal.
2Zn(s)+O2(g)→2ZnO(s)
Express your answer with the appropriate units.
1
Expert's answer
2020-11-23T06:56:58-0500

"n = m\/M"

"M(Zn) = 65g\/mol"

From the equation it follows that "n(Zn)= n(ZnO) = 3/65\ mol"

"m = n*M"

"M(ZnO)= 81g\/mol"

"m(ZnO) = 81 * 3/65 g \approx 3.7385g"


Answer: "m(ZnO)= 3.7385g"